# Dependability of UAV-Based Networks and Computing Systems: A Survey

### v1

@@ 6.1. Fault Prevention
Fault Prevention | Energy & Resource-Aware Design | (Ramos-Ramos et al., 2024), (Zhou et al., 2022a), (Wei et al., 2022), (Jeong et al., 2020b), (Zhang and Bai, 2024), (Fatima et al., 2024a), (Maatouk et al., 2024), (Xu et al., 2020), (Suganya et al., 2024), (Peng et al., 2024), (Palossi et al., 2019), (Rahul et al., 2021), (Wang et al., 2022b), (Asim et al., 2022), (Zhang et al., 2023a), (Alioua et al., 2018), (Zhang et al., 2025), (Yang LI, 2024), (Selim et al., 2021), (Xue et al., 2023a), (Tang et al., 2024a), (Wang et al., 2020e), (Laroui et al., 2021), (Lin et al., 2021a), (Zhang et al., 2021b), (Raj and Simmhan, 2023), (Shao et al., 2023b), (Seisa et al., 2022b), (Hossain et al., 2017), (Lu et al., 2024), (Grasso and Schembra, 2019), (Dou et al., 2024), (Ma and Sun, 2022a), (Qureshi et al., 2024a), (Grasso et al., 2022a), (Gupta et al., 2021), (Wang et al., 2019a), (Chen et al., 2020a), (Awada et al., 2023a), (Navardi and Mohsenin, 2023b), (Martínez-Sánchez et al., 2023), (Vélez et al., 2024), (Faraci et al., 2023a), (Zeng et al., 2024b), (Shi et al., 2023), (Hong et al., 2022), (Grasso et al., 2022b), (Baktayan et al., 2022), (Hang et al., 2020a), (Tsung et al., 2022), (Cheng et al., 2023), (Asheralieva and Niyato, 2020), (Geng et al., 2024a), (Shams et al., 2023), (Gao et al., 2019a), (Akbari et al., 2024a), (Li et al., 2024a), (Hadjkouider et al., 2023), (Zheng et al., 2024a), (Karam et al., 2024a), (Xiang et al., 2024), (Venkatagiri et al., 2018b), (Liu et al., 2019), (Callegaro et al., 2021), (Khan et al., 2022), (Li et al., 2020c), (Zhao et al., 2022), (Xu et al., 2024a), (Asheralieva and Niyato, 2019), (Yu et al., 2022a), (Huang et al., 2024b), (Apostolopoulos et al., 2023), (Jung et al., 2023b), (Zhang et al., 2018a), (Wang et al., 2023b), (Haber et al., 2021a), (Pandey et al., 2021a), (Wang et al., 2024b), (Dai et al., 2024), (Alam et al., 2019), (Subburaj et al., 2023), (Fatima et al., 2024b), (Yang et al., 2020), (Gu et al., 2022a), (Liu et al., 2023c), (Falcão et al., 2024), (Wang et al., 2020d), (Hou et al., 2020d), (Lu et al., 2022), (Yu et al., 2021), (Kim and Kim, 2021), (Manzoor et al., 2021), (Zhang et al., 2024d), (Thapliyal et al., 2024a), (Awada et al., 2022a), (Li et al., 2020a), (Li et al., 2024f), (Tang et al., 2024c), (Luo et al., 2023), (Lin et al., 2022), (Ghosh and Kuila, 2023), (Islam et al., 2019), (Ye et al., 2020), (Yu et al., 2022b), (Ju et al., 2024a), (Ouyang et al., 2023), (Rao and Naidu, 2024), (Awada et al., 2023b), (Zhang et al., 2024b), (Rokaha et al., 2019), (Hao et al., 2024), (Tang et al., 2022a), (Zhang et al., 2023e), (Pan et al., 2022), (Xu et al., 2022), (Liu et al., 2023b), (Haber et al., 2019b), (Wang et al., 2024a), (Liang et al., 2021), (Tang et al., 2024b), (Zeng et al., 2023), (Wang et al., 2021b), (Hou et al., 2020b), (Chu et al., 2021a), (Tang et al., 2022b), (Ju et al., 2024b), (Chen et al., 2020c), (Yuan et al., 2022), (Falcão et al., 2023a), (Liu et al., 2024b), (Hossain et al., 2022), (Chen et al., 2020b), (Xue et al., 2023c), (Hadi and Ghazizadeh, 2023), (Gao et al., 2020), (Kumar and Jannesari, 2023), (Pandey et al., 2021b), (Jeong et al., 2020a), (He et al., 2023b), (Han et al., 2021), (Dai et al., 2022), (Zhang et al., 2024a), (Wang et al., 2021g), (Zhang et al., 2022b), (Huang and Wang, 2024), (Zhang et al., 2022a), (Wei et al., 2021), (Qu et al., 2022), (Chai et al., 2023), (Wang et al., 2018), (Zhang et al., 2023f), (Zhang et al., 2024e), (Asim and Abd El-Latif, 2023), (Ouyang et al., 2024), (Seong et al., 2024), (Yu et al., 2023b), (Awada et al., 2022b), (Geng et al., 2024b), (Awada et al., 2021a), (Zeng et al., 2024c), (Zhang et al., 2023d), (Hou et al., 2020a), (Khoramnejad et al., 2024), (Chen et al., 2024a), (Wang et al., 2020c), (Zheng et al., 2023), (Chu et al., 2021b), (Li et al., 2024a), (Qureshi et al., 2024b), (Hou et al., 2020c), (Wang et al., 2024c), (Tian et al., 2024), (Zhang et al., 2018b), (Bezziane et al., 2024), (Chen et al., 2021a), (Basharat et al., 2023), (Chen et al., 2022b), (Zou et al., 2022), (D’Agati et al., 2024), (Zhao et al., 2024), (Su et al., 2022), (Thapliyal et al., 2024b), (Gao et al., 2019b), (Ma and Sun, 2022b), (Akbari et al., 2024b), (Xiong et al., 2019), (Al-habob et al., 2024), (Sun et al., 2020), (Navardi and Mohsenin, 2023a), (Faraci et al., 2023b), (Seisa et al., 2022a), (Luo et al., 2021a), (Haber et al., 2019a), (Bao et al., 2023), (Bai et al., 2023c), (Jung et al., 2023a), (Ngoenriang et al., 2022a), (Kong et al., 2022c), (Ning et al., 2024), (He et al., 2023a), (Gupta and Gupta, 2022), (Picano and Fantacci, 2024b), (Wang et al., 2021a), (Moeyersons et al., 2021), (Kong et al., 2022b), (Shao et al., 2023a), (Pamuklu et al., 2023), (Kaleem et al., 2019a), (Lin et al., 2021b), (Garge and Balakrishna, 2018), (Hang et al., 2020b), (Tendolkar et al., 2021), (Hui et al., 2024), (Xue et al., 2023b), (Gao et al., 2024), (Manzini et al., 2023a), (Liu et al., 2024c), (Ju et al., 2023), (Pang et al., 2024) |
@@
Robust Trajectory & Collision-Avoidance Planning | (Lee et al., 2022b), (Krishnan et al., 2021), (Alqudsi et al., 2023), (Wang et al., 2022b), (Hu et al., 2024), (Asim et al., 2022), (Zhang et al., 2023a), (Xu et al., 2024c), (Singh et al., 2023), (Souli et al., 2023), (Yang LI, 2024), (Li et al., 2024c), (Kaymaz et al., 2024), (Zeng et al., 2024a), (Hirota et al., 2020b), (Gao et al., 2021a), (Chen et al., 2022a), (Zheng et al., 2024a), (Liu et al., 2019), (Huang et al., 2024b), (Wang et al., 2023b), (Choutri et al., 2022), (Dai et al., 2024), (Subburaj et al., 2023), (Liu et al., 2023c), (Wang et al., 2020d), (Hou et al., 2020d), (Lu et al., 2022), (Yu et al., 2021), (Luo et al., 2021b), (Tripathi et al., 2023a), (Tang et al., 2024c), (Lee et al., 2022a), (Luo et al., 2023), (Dai et al., 2018), (Ye et al., 2020), (Behjat et al., 2019), (Koerkamp et al., 2019), (She and Ouyang, 2021), (Huang et al., 2019), (Ju et al., 2024a), (Awada et al., 2023b), (Li et al., 2018), (Deng et al., 2022), (Zhou et al., 2024a), (Baccarelli et al., 2024), (Li et al., 2024d), (Kim et al., 2024a), (Shen et al., 2023b), (Huang et al., 2024a), (Xue et al., 2023c), (Peng et al., 2022), (Lv et al., 2024), (Qayyum et al., 2021), (Yang et al., 2022), (Lorincz et al., 2021a), (Wang et al., 2021g), (Yang et al., 2019) (Lindqvist et al., 2020), (Ye et al., 2022), (Guo et al., 2022a), (Bekkouche et al., 2018), (Tripathi et al., 2023b), (De Paula Soares et al., 2023), (Zhan and Xu, 2024), (Zhou et al., 2022b), (Li et al., 2024e), (Gao et al., 2021b), (Hirota et al., 2020a), (Wu et al., 2023), (Zheng et al., 2024b), (Jiang et al., 2019), (Ngoenriang et al., 2022a), (Kong et al., 2022c), (Ning et al., 2024), (Wang et al., 2024e), (Marcos and Utete, 2021) |
@@ 6.1.1. Energy & Resource-Aware Design:
For instance, energy-aware offloading frameworks are proposed to reduce UAV energy consumption by dynamically assigning tasks to ground servers or other UAVs based on current battery levels and computational demands (Ramos-Ramos et al., 2024; Zhou et al., 2022a; Fatima et al., 2024a; Wang et al., 2022b; Yang LI, 2024; Suganya et al., 2024; Peng et al., 2024; Zhang et al., 2025; Wang et al., 2020e; Seisa et al., 2022b; Chen et al., 2020a; Faraci et al., 2023a; Asheralieva and Niyato, 2020; Li et al., 2024a; Liu et al., 2019; Callegaro et al., 2021; Jung et al., 2023b; Haber et al., 2021a; Yang et al., 2020; Awada et al., 2022a). Additionally, many works integrate power control and bandwidth allocation schemes to adaptively regulate UAV transmissions and processing loads, thereby lowering overall energy usage while maintaining service quality (Zhang and Bai, 2024; Xu et al., 2020; Selim et al., 2021; Tang et al., 2024a; Hossain et al., 2017; Hang et al., 2020a; Cheng et al., 2023; Gao et al., 2019a; Xiang et al., 2024; Zhao et al., 2022; Zhang et al., 2018a; Wang et al., 2023b; Liu et al., 2023c; Kim and Kim, 2021). These strategies directly contribute to enhancing system availability and reliability by preventing unexpected shutdowns and maintaining functional operation under varying load conditions.
@@ 6.1.2. Robust Trajectory & Collision-Avoidance Planning:
Multiple studies have proposed robust trajectory planning and avoidance frameworks that balance safety with performance. For example, model predictive control, velocity obstacles, and tabu search methods are commonly used to generate safe flight paths in constrained environments (Lee et al., 2022b; Wang et al., 2022b; Xu et al., 2024c; Bekkouche et al., 2018; Wu et al., 2023). Reinforcement learning and DRL-based approaches further enhance adaptivity, allowing UAVs to dynamically adjust paths in response to moving obstacles or real-time conditions (Hu et al., 2024; Behjat et al., 2019; Gao et al., 2021b; Kong et al., 2022c; Ning et al., 2024). Some works also incorporate probabilistic models or risk-aware strategies to handle uncertainties such as wind gusts, localization errors, or multi-agent interactions (Zeng et al., 2024a; Liu et al., 2019; Luo et al., 2021b; Zhan and Xu, 2024).
@@ 6.2. Fault tolerance
Fault Tolerance | Error Detection | (Ahamad and Mishra, 2023), (Cheng et al., 2024), (Vemulapalli and Mesapam, 2021), (Gokulakrishnan et al., 2023), (Ma et al., 2024), (Saadaoui et al., 2023), (Kostoglotov et al., 2024), (Zhang and Li, 2018), (Zhou et al., 2024d), (Amin and Obermaisser, 2018b), (Jeong et al., 2017), (Wang et al., 2021c), (Mo et al., 2022), (Chang et al., 2024), (Alam et al., 2019), (Wu et al., 2024), (Balamuralidhar et al., 2021), (Behjat et al., 2019), (Bakhtiari et al., 2020b), (Rokaha et al., 2019), (Nwadiugwu et al., 2021), (Zhong et al., 2021), (Jeong et al., 2016), (Shang et al., 2023), (Pan et al., 2024), (Li et al., 2024b), (Zhou et al., 2023), (Zhang and Zhang, 2024), (Zheng et al., 2021), (Yuwei and Aijun, 2020) |
@@
Recovery | (Farrukh and West, 2023), (Luo et al., 2022), (Bobronnikov et al., 2022), (Alkinani et al., 2021), (Zuo et al., 2022), (Wang et al., 2021f), (Zhou and Cao, 2018), (Callegaro et al., 2021), (Choutri et al., 2022), (Wang et al., 2024b), (Hou et al., 2020d), (Luo et al., 2021b), (Machida and Andrade, 2021a), (Kim and Kim, 2021), (Li et al., 2024f), (Machida et al., 2023a), (Yu et al., 2022b), (Ju et al., 2024a), (Awada et al., 2023b), (Zhang et al., 2024b), (Machida et al., 2023b), (Petrenko et al., 2023), (Pan et al., 2022), (Watanabe and Machida, 2022), (Wang et al., 2019c), (Li et al., 2023a), (Damigos et al., 2023), (Machida and Andrade, 2021b) | |
@@
Redundant Architecture | (Ma et al., 2024), (Luo et al., 2020), (Farrukh and West, 2023), (La Salandra et al., 2024), (Ren et al., 2017), (Kilbourne et al., 2021), (Nguyen et al., 2023), (Patnayak et al., 2021a), (Awada et al., 2021b), (Zhang et al., 2021c), (Gallego-Madrid et al., 2020), (Kaleem et al., 2019b), (Bian et al., 2023), (Dogea et al., 2023), (Karam et al., 2024a), (Aloqaily et al., 2021), (Liu et al., 2019), (Callegaro et al., 2021), (Khan et al., 2022), (Zhao et al., 2022), (Jung et al., 2023b), (Sharma et al., 2019a), (Pacheco et al., 2021b), (Hou et al., 2020d), (Awada et al., 2022a), (Ali and Ali, 2022), (Kim, 2024), (Petrenko et al., 2023), (Sun et al., 2018), (El Sayed et al., 2023), (Sara et al., 2016), (Chen and Ma, 2024), (Barnawi et al., 2021), (Wang et al., 2020a), (Javanmardi et al., 2024), (Panwar et al., 2023), (Shen et al., 2023a), (Naouri et al., 2024), (Tang et al., 2023), (Pustokhina et al., 2021), (Taleb et al., 2023a), (Patnayak et al., 2021b), (Lorincz et al., 2021b), (Karam et al., 2024b), (Lei et al., 2023), (Li et al., 2023b), (Haber et al., 2021b), (Li and Nie, 2023), (Pradhan et al., 2015) | |
@@
Network Topology Design | (Xu et al., 2021), (Guruprasad et al., 2024), (Zhang et al., 2023b), (Vamvakas et al., 2019), (Wang et al., 2024d), (Chen et al., 2022c), (Fujii et al., 2017), (Shao et al., 2023b), (Martínez-Sánchez et al., 2023), (Wang et al., 2021e), (Dao et al., 2021), (Ozger et al., 2018b), (A. Alissa et al., 2023), (Tan et al., 2020), (Su, 2021), (Khan et al., 2021), (Wang et al., 2022c), (Zhou et al., 2024b), (Kong et al., 2022a), (Chen et al., 2024b), (Aloqaily et al., 2021), (Khan et al., 2022), (Li et al., 2020c), (Haber et al., 2021a), (Sharma et al., 2019a), (Gu et al., 2022a), (Thapliyal et al., 2024a), (Li et al., 2020b), (Zhang et al., 2021a), (Jung et al., 2023c), (Mukherjee et al., 2020), (Paranjothi et al., 2019), (Peng et al., 2023), (Nwadiugwu et al., 2021), (Mukherjee et al., 2023), (Mao et al., 2023), (Zhang et al., 2023c), (Yue et al., 2018), (Yadav and Baranwal, 2023), (Wang et al., 2020b), (Zhang et al., 2024c), (Amos et al., 2021), (Khanna et al., 2019), (Wang et al., 2023a), (Wang et al., 2022d), (Zhou et al., 2024c), (Liu et al., 2024d), (Liu et al., 2023a), (Taleb et al., 2023b), (Gu et al., 2022b), (Wang et al., 2019b), (Sharma et al., 2019b), (Ren et al., 2024), (Shao et al., 2023a), (Song et al., 2022), (Li et al., 2020d), (Ozger et al., 2018a) |

### v2

@@ 6.1. Fault Prevention
| Fault Prevention | Energy & Resource-Aware Design | (Ramos-Ramos et al., 2024), (Zhou et al., 2022a), (Wei et al., 2022), (Jeong et al., 2020), (Zhang and Bai, 2024), (Fatima et al., 2024), (Maatouk et al., 2024), (Xu et al., 2020), (Suganya et al., 2024), (Peng et al., 2024), (Palossi et al., 2019), (Rahul et al., 2021), (Wang et al., 2022b), (Asim et al., 2022), (Zhang et al., 2023a), (Alioua et al., 2018), (Zhang et al., 2025), (Yang LI, 2024), (Selim et al., 2021), (Xue et al., 2023a), (Tang et al., 2024a), (Wang et al., 2020d), (Laroui et al., 2021), (Lin et al., 2021), (Zhang et al., 2021b), (Raj and Simmhan, 2023), (Shao et al., 2023), (Seisa et al., 2022), (Hossain et al., 2017), (Lu et al., 2024), (Grasso and Schembra, 2019), (Dou et al., 2024), (Ma and Sun, 2022), (Qureshi et al., 2024), (Grasso et al., 2022a), (Gupta et al., 2021), (Wang et al., 2019a), (Chen et al., 2020a), (Awada et al., 2023), (Navardi and Mohsenin, 2023), (Mart´ınez-Sánchez et al., 2023), (Vélez et al., 2024), (Faraci et al., 2023), (Zeng et al., 2024b), (Shi et al., 2023), (Hong et al., 2022), (Grasso et al., 2022b), (Baktayan et al., 2022), (Hang et al., 2020), (Tsung et al., 2022), (Cheng et al., 2023), (Asheralieva and Niyato, 2020), (Geng et al., 2024), (Shams et al., 2023), (Gao et al., 2019), (Akbari et al., 2024), (Li et al., 2024a), (Hadjkouider et al., 2023), (Zheng et al., 2024), (Karam et al., 2024), (Xiang et al., 2024), (Venkatagiri et al., 2018), (Liu et al., 2019), (Callegaro et al., 2021), (Khan et al., 2022), (Li et al., 2020c), (Zhao et al., 2022), (Xu et al., 2024a), (Asheralieva and Niyato, 2019), (Yu et al., 2022a), (Huang et al., 2024b), (Apostolopoulos et al., 2023), (Jung et al., 2023a), (Zhang et al., 2018), (Wang et al., 2023b), (Haber et al., 2021), (Pandey et al., 2021a), (Wang et al., 2024b), (Dai et al., 2024), (Alam et al., 2019), (Subburaj et al., 2023), (Fatima et al., 2024), (Yang et al., 2020), (Gu et al., 2022), (Liu et al., 2023c), (Falcão et al., 2024), (Wang et al., 2020c), (Hou et al., 2020b), (Lu et al., 2022), (Yu et al., 2021), (Kim and Kim, 2021), (Manzoor et al., 2021), (Zhang et al., 2024d), (Thapliyal et al., 2024), (Awada et al., 2022), (Li et al., 2020a), (Li et al., 2024e), (Tang et al., 2024c), (Luo et al., 2023), (Lin et al., 2022), (Ghosh and Kuila, 2023), (Islam et al., 2019), (Ye et al., 2020), (Yu et al., 2022b), (Ju et al., 2024), (Ouyang et al., 2023), (Rao and Naidu, 2024), (Awada et al., 2023), (Zhang et al., 2024b), (Rokaha et al., 2019), (Hao et al., 2024), (Tang et al., 2022a), (Zhang et al., 2023d), (Pan et al., 2022), (Xu et al., 2022), (Liu et al., 2023b), (Haber et al., 2019), (Wang et al., 2024a), (Liang et al., 2021), (Tang et al., 2024b), (Zeng et al., 2023), (Wang et al., 2021a), (Hou et al., 2020a), (Chu et al., 2021), (Tang et al., 2022b), (Ju et al., 2024), (Chen et al., 2020c), (Yuan et al., 2022), (Falcão et al., 2023), (Liu et al., 2024b), (Hossain et al., 2022), (Chen et al., 2020b), (Xue et al., 2023c), (Hadi and Ghazizadeh, 2023), (Gao et al., 2020), (Kumar and Jannesari, 2023), (Pandey et al., 2021b), (Jeong et al., 2020), (He et al., 2023), (Han et al., 2021), (Dai et al., 2022), (Zhang et al., 2024a), (Wang et al., 2021e), (Zhang et al., 2022b), (Huang and Wang, 2024), (Zhang et al., 2022a), (Wei et al., 2021), (Qu et al., 2022), (Chai et al., 2023), (Wang et al., 2018), (Zhang et al., 2023e), (Zhang et al., 2024e), (Asim and Abd El-Latif, 2023), (Ouyang et al., 2024), (Seong et al., 2024), (Yu et al., 2023b), (Awada et al., 2022), (Geng et al., 2024), (Awada et al., 2021), (Zeng et al., 2024b), (Zhang et al., 2023c), (Hou et al., 2020a), (Khoramnejad et al., 2024), (Chen et al., 2024a), (Wang et al., 2020c), (Zheng et al., 2023), (Chu et al., 2021), (Li et al., 2024a), (Qureshi et al., 2024), (Hou et al., 2020b), (Wang et al., 2024b), (Tian et al., 2024), (Zhang et al., 2018), (Bezziane et al., 2024), (Chen et al., 2021a), (Basharat et al., 2023), (Chen et al., 2022a), (Zou et al., 2022), (D’Agati et al., 2024), (Zhao et al., 2024), (Su et al., 2022), (Thapliyal et al., 2024), (Gao et al., 2019), (Ma and Sun, 2022), (Akbari et al., 2024), (Xiong et al., 2019), (Al-habob et al., 2024), (Sun et al., 2020), (Navardi and Mohsenin, 2023), (Faraci et al., 2023), (Seisa et al., 2022), (Luo et al., 2021), (Haber et al., 2019), (Bao et al., 2023), (Bai et al., 2023b), (Jung et al., 2023a), (Ngoenriang et al., 2022), (Kong et al., 2022c), (Ning et al., 2024), (He et al., 2023), (Gupta and Gupta, 2022), (Picano and Fantacci, 2024), (Wang et al., 2021a), (Moeyersons et al., 2021), (Kong et al., 2022b), (Shao et al., 2023), (Pamuklu et al., 2023), (Kaleem et al., 2019), (Lin et al., 2021), (Garge and Balakrishna, 2018), (Hang et al., 2020), (Tendolkar et al., 2021), (Hui et al., 2024), (Xue et al., 2023b), (Gao et al., 2024), (Manzini et al., 2023), (Liu et al., 2024b), (Ju et al., 2023), (Pang et al., 2024) |
@@
| Robust Trajectory & Collision-Avoidance Planning | (Lee et al., 2022b), (Krishnan et al., 2021), (Alqudsi et al., 2023), (Wang et al., 2022b), (Hu et al., 2024), (Asim et al., 2022), (Zhang et al., 2023a), (Xu et al., 2024c), (Singh et al., 2023), (Souli et al., 2023), (Yang LI, 2024), (Li et al., 2024c), (Kaymaz et al., 2024), (Zeng et al., 2024a), (Hirota et al., 2020), (Gao et al., 2021), (Chen et al., 2022a), (Zheng et al., 2024), (Liu et al., 2019), (Huang et al., 2024b), (Wang et al., 2023b), (Choutri et al., 2022), (Dai et al., 2024), (Subburaj et al., 2023), (Liu et al., 2023c), (Wang et al., 2020c), (Hou et al., 2020b), (Lu et al., 2022), (Yu et al., 2021), (Luo et al., 2021), (Tripathi et al., 2023), (Tang et al., 2024c), (Lee et al., 2022a), (Luo et al., 2023), (Dai et al., 2018), (Ye et al., 2020), (Behjat et al., 2019), (Koerkamp et al., 2019), (She and Ouyang, 2021), (Huang et al., 2019), (Ju et al., 2024), (Awada et al., 2023), (Li et al., 2018), (Deng et al., 2022), (Zhou et al., 2024a), (Baccarelli et al., 2024), (Li et al., 2024d), (Kim et al., 2024a), (Shen et al., 2023b), (Huang et al., 2024a), (Xue et al., 2023c), (Peng et al., 2022), (Lv et al., 2024), (Qayyum et al., 2021), (Yang et al., 2022), (Lorincz et al., 2021), (Wang et al., 2021e), (Yang et al., 2019) (Lindqvist et al., 2020), (Ye et al., 2022), (Guo et al., 2022a), (Bekkouche et al., 2018), (Tripathi et al., 2023), (De Paula Soares et al., 2023), (Zhan and Xu, 2024), (Zhou et al., 2022b), (Li et al., 2024e), (Gao et al., 2021), (Hirota et al., 2020), (Wu et al., 2023), (Zheng et al., 2024), (Jiang et al., 2019), (Ngoenriang et al., 2022), (Kong et al., 2022c), (Ning et al., 2024), (Wang et al., 2024d), (Marcos and Utete, 2021) |
@@ 6.1.1. Energy & Resource-Aware Design:
For instance, energy-aware offloading frameworks are proposed to reduce UAV energy consumption by dynamically assigning tasks to ground servers or other UAVs based on current battery levels and computational demands (Ramos-Ramos et al., 2024; Zhou et al., 2022a; Fatima et al., 2024; Wang et al., 2022b; Yang LI, 2024; Suganya et al., 2024; Peng et al., 2024; Zhang et al., 2025; Wang et al., 2020d; Seisa et al., 2022; Chen et al., 2020a; Faraci et al., 2023; Asheralieva and Niyato, 2020; Li et al., 2024a; Liu et al., 2019; Callegaro et al., 2021; Jung et al., 2023a; Haber et al., 2021; Yang et al., 2020; Awada et al., 2022). Additionally, many works integrate power control and bandwidth allocation schemes to adaptively regulate UAV transmissions and processing loads, thereby lowering overall energy usage while maintaining service quality (Zhang and Bai, 2024; Xu et al., 2020; Selim et al., 2021; Tang et al., 2024a; Hossain et al., 2017; Hang et al., 2020; Cheng et al., 2023; Gao et al., 2019; Xiang et al., 2024; Zhao et al., 2022; Zhang et al., 2018; Wang et al., 2023b; Liu et al., 2023c; Kim and Kim, 2021). These strategies directly contribute to enhancing system availability and reliability by preventing unexpected shutdowns and maintaining functional operation under varying load conditions.
@@ 6.1.2. Robust Trajectory & Collision-Avoidance Planning:
Multiple studies have proposed robust trajectory planning and avoidance frameworks that balance safety with performance. For example, model predictive control, velocity obstacles, and tabu search methods are commonly used to generate safe flight paths in constrained environments (Lee et al., 2022b; Wang et al., 2022b; Xu et al., 2024c; Bekkouche et al., 2018; Wu et al., 2023). Reinforcement learning and DRL-based approaches further enhance adaptivity, allowing UAVs to dynamically adjust paths in response to moving obstacles or real-time conditions (Hu et al., 2024; Behjat et al., 2019; Gao et al., 2021; Kong et al., 2022c; Ning et al., 2024). Some works also incorporate probabilistic models or risk-aware strategies to handle uncertainties such as wind gusts, localization errors, or multi-agent interactions (Zeng et al., 2024a; Liu et al., 2019; Luo et al., 2021; Zhan and Xu, 2024).
@@ 6.2. Fault tolerance
| Fault Tolerance | Error Detection | (Ahamad and Mishra, 2023), (Cheng et al., 2024), (Vemulapalli and Mesapam, 2021), (Gokulakrishnan et al., 2023), (Ma et al., 2024), (Saadaoui et al., 2023), (Kostoglotov et al., 2024), (Zhang and Li, 2018), (Zhou et al., 2024c), (Amin and Obermaisser, 2018), (Jeong et al., 2017), (Wang et al., 2021b), (Mo et al., 2022), (Chang et al., 2024), (Alam et al., 2019), (Wu et al., 2024), (Balamuralidhar et al., 2021), (Behjat et al., 2019), (Bakhtiari et al., 2020), (Rokaha et al., 2019), (Nwadiugwu et al., 2021), (Zhong et al., 2021), (Jeong et al., 2016), (Shang et al., 2023), (Pan et al., 2024), (Li et al., 2024b), (Zhou et al., 2023), (Zhang and Zhang, 2024), (Zheng et al., 2021), (Yuwei and Aijun, 2020) |
@@
| Recovery | (Farrukh and West, 2023), (Luo et al., 2022), (Bobronnikov et al., 2022), (Alkinani et al., 2021), (Zuo et al., 2022), (Wang et al., 2021d), (Zhou and Cao, 2018), (Callegaro et al., 2021), (Choutri et al., 2022), (Wang et al., 2024b), (Hou et al., 2020b), (Luo et al., 2021), (Machida and Andrade, 2021), (Kim and Kim, 2021), (Li et al., 2024e), (Machida et al., 2023), (Yu et al., 2022b), (Ju et al., 2024), (Awada et al., 2023), (Zhang et al., 2024b), (Machida et al., 2023), (Petrenko et al., 2023), (Pan et al., 2022), (Watanabe and Machida, 2022), (Wang et al., 2019b), (Li et al., 2023a), (Damigos et al., 2023), (Machida and Andrade, 2021) | |
@@
| Redundant Architecture | (Ma et al., 2024), (Luo et al., 2020), (Farrukh and West, 2023), (La Salandra et al., 2024), (Ren et al., 2017), (Kilbourne et al., 2021), (Nguyen et al., 2023), (Patnayak et al., 2021), (Awada et al., 2021), (Zhang et al., 2021c), (Gallego-Madrid et al., 2020), (Kaleem et al., 2019), (Bian et al., 2023), (Dogea et al., 2023), (Karam et al., 2024), (Aloqaily et al., 2021), (Liu et al., 2019), (Callegaro et al., 2021), (Khan et al., 2022), (Zhao et al., 2022), (Jung et al., 2023a), (Sharma et al., 2019), (Pacheco et al., 2021), (Hou et al., 2020b), (Awada et al., 2022), (Ali and Ali, 2022), (Kim, 2024), (Petrenko et al., 2023), (Sun et al., 2018), (El Sayed et al., 2023), (Sara et al., 2016), (Chen and Ma, 2024), (Barnawi et al., 2021), (Wang et al., 2020a), (Javanmardi et al., 2024), (Panwar et al., 2023), (Shen et al., 2023a), (Naouri et al., 2024), (Tang et al., 2023), (Pustokhina et al., 2021), (Taleb et al., 2023), (Patnayak et al., 2021), (Lorincz et al., 2021), (Karam et al., 2024), (Lei et al., 2023), (Li et al., 2023b), (Haber et al., 2021), (Li and Nie, 2023), (Pradhan et al., 2015) | |
@@
| Network Topology Design | (Xu et al., 2021), (Guruprasad et al., 2024), (Zhang et al., 2023b), (Vamvakas et al., 2019), (Wang et al., 2024c), (Chen et al., 2022b), (Fujii et al., 2017), (Shao et al., 2023), (Mart´ınez-Sánchez et al., 2023), (Wang et al., 2021c), (Dao et al., 2021), (Ozger et al., 2018), (A. Alissa et al., 2023), (Tan et al., 2020), (Su, 2021), (Khan et al., 2021), (Wang et al., 2022c), (Zhou et al., 2024b), (Kong et al., 2022a), (Chen et al., 2024b), (Aloqaily et al., 2021), (Khan et al., 2022), (Li et al., 2020c), (Haber et al., 2021), (Sharma et al., 2019), (Gu et al., 2022), (Thapliyal et al., 2024), (Li et al., 2020b), (Zhang et al., 2021a), (Jung et al., 2023b), (Mukherjee et al., 2020), (Paranjothi et al., 2019), (Peng et al., 2023), (Nwadiugwu et al., 2021), (Mukherjee et al., 2023), (Mao et al., 2023), (Zhang et al., 2023b), (Yue et al., 2018), (Yadav and Baranwal, 2023), (Wang et al., 2020b), (Zhang et al., 2024c), (Amos et al., 2021), (Khanna et al., 2019), (Wang et al., 2023a), (Wang et al., 2022c), (Zhou et al., 2024b), (Liu et al., 2024c), (Liu et al., 2023a), (Taleb et al., 2023), (Gu et al., 2022), (Wang et al., 2019a), (Sharma et al., 2019), (Ren et al., 2024), (Shao et al., 2023), (Song et al., 2022), (Li et al., 2020c), (Ozger et al., 2018) |